a)     A simple closed economy with an mpc equal to 0.5. Investment spending has suddenly fallen, reducing aggregate demand and output to a level that is 100 million below Y*.

                  

               iii.           If the government decided to try to get the economy back to full employment using only an increase in transfers, how large would this increase need to be?


1
Expert's answer
2021-07-05T16:16:49-0400

For a simple closed 3-sector economy like the above, the following Keynesian framework can be used to represent it.


"Y = C + I + G,"


where Y = national income, C = consumption expenditure, I = investment expenditure and G = government expenditure.


Now,"C = a + bY' and Y' = Y - T + R,"


where a = autonomous consumption, b = marginal propensity to consume, Y' = disposable income, T = taxes, and R = transfers.


So we have "C = a + b(Y - T + R) = a + bY - bT + bR."


Substituting C in the original national income equation.


"Y = a + bY - bT + bR + I + G,"


or


"Y - bY = a - bT + bR + I + G,"


or


"Y = \frac{1}{(1-b)} (a - bT + bR + I + G) \Rightarrow \textup{ Equation 1}"


Using this equation we can calculate the impacts of changes in various variables on the changes in national income.


To see how Y changes with respect to change in G or government spending, we will simply differentiate Equation 1 wrt G; such as


"\frac{\mathrm{d} Y}{\mathrm{d} G} = \frac{\mathrm{d} (\frac{1}{(1-b)} (a - bT + bR + I + G)) }{\mathrm{d} G} = \frac{1}{(1-b)}"


Thus for a unit change in G, Y changes by 1/(1-b), where b is mpc, like stated previously. This is the government expenditure or spending multiplier. We can state that,


"\textup{Government Spending Multiplier} = \textup{GM} = \frac{\mathrm{d} Y}{\mathrm{d} G} = \frac{1}{(1-b)}"


Similarly, we can calculate the tax multiplier and transfers multiplier.


Tax multiplier:


"\frac{\mathrm{d} Y}{\mathrm{d} T} = \frac{\mathrm{d} (\frac{1}{(1-b)} (a - bT + bR + I + G)) }{\mathrm{d} T} = \frac{-b}{(1-b)}" }


"\textup{Tax Multiplier} = \textup{TM} = \frac{\mathrm{d} Y}{\mathrm{d} T} = \frac{-b}{(1-b)}"


Transfers multiplier:


"\frac{\mathrm{d} Y}{\mathrm{d} R} = \frac{\mathrm{d} (\frac{1}{(1-b)} (a - bT + bR + I + G)) }{\mathrm{d} R} = \frac{b}{(1-b)}"


"\textup{Transfers Multiplier} = \textup{RM} = \frac{\mathrm{d} Y}{\mathrm{d} R} = \frac{b}{(1-b)}"


Now we are given that mpc = b = 0.5. So


"\textup{GM} = \frac{\mathrm{d} Y}{\mathrm{d} G} = \frac{1}{(1-b)} = \frac{1}{(1-0.5)} = 2"


"\textup{TM} = \frac{\mathrm{d} Y}{\mathrm{d} T} = \frac{-b}{(1-b)} = \frac{-0.5}{(1-0.5)} = -1"


"\textup{RM} = \frac{\mathrm{d} Y}{\mathrm{d} R} = \frac{b}{(1-b)} = \frac{0.5}{(1-0.5)} = 1"


Now the needed change in national income is 100 million, so as to bring it to the level of full employment equilibrium. We are given three cases. Let's approach each individually.


Case 1: When G has to vary


Earlier we have calculated GM = 2. And GM = change in Y / change in G = 2. When change in Y = 100, change in G = 50. Thus, government spending must increase by 50 million, to increase national income by 100 million.


Case 2: When T has to vary


TM = change in Y / change in T = -1. When change in Y = 100, change in T = -100. Thus, taxes must be decreased by 100 million, to increase national income by 100 million.


Case 3: When R has to vary


RM = change in Y / change in R = 1. When change in Y = 100, change in R = 100. Thus, transfers must be increased by 100 million, to increase national income by 100 million.